OpenStudy (anonymous):

Triangle ABC has side lengths 7, 24, and 25. Do the side lengths form a Pythagorean triple? Please explain it to me.

OpenStudy (rogue):

The sides form a Pythagorean triple if they satisfy the Pythagorean theorem, which is\[a^2 + b^2 = c^2\]In your case, a = 7, b = 24, c = 25.\[7^2 + 24^2 = 25^2\]\[49 + 576 = 625\]\[625 = 625\]It satisfies the Pythagorean theorem, and so the side lengths are a triple.

OpenStudy (anonymous):

7^2 = 49, 25^2 = 625, 24^2 = 576 Since 576 + 49 = 625 a^2 + b^2 = c^2 Yes, it's right triangle!

OpenStudy (anonymous):

Thank you so much, I was having so much trouble before, thank you for explaining.